# PRIMA.CPP: Speeding Up 70B-Scale LLM Inference on Low-Resource Everyday Home Clusters

### v1

@@ 3.2 Layer-to-device Assignment Problem
An implicit assumption was made in section 3.1: all devices, even heterogeneous ones, use the same layer window size, which is poor practice. High-end devices should handle more layers than low-end ones. As shown in Figure 6(f), if we offload some layers from devices 2, 3 to devices 1, 4, “bubbles" (device idle time) can be smaller, and more rounds can be executed within the same time. So here comes the question: how should the model layers be distributed among resource-limited and heterogeneous devices, and how many layers should run on the GPU (if any)? To answer this question, we define the layer-to-device assignment (LDA) problem as follows.
Definition 1 (Layer-to-Device Assignment Problem, LDA).
Assume there are $M$ devices, $w_{m}$ is the layer window size on device $d_{m}$ and $n_{m}$ is the number of GPU layers within $w_{m}$. Let the decision variables be ${\bm{w}}^{\mathrm{T}}=[w_{1},w_{2},\cdots,w_{M}]$ and ${\bm{n}}^{\mathrm{T}}=[n_{1},n_{2},\cdots,n_{M}]$. Our objective is to find the optimal ${\bm{w}}$ and ${\bm{n}}$ that minimizes the token latency:
@@
where $L$ is the number of model layers; ${\bm{a}},{\bm{b}},{\bm{c}}$ are constant vectors determined by computation latency, memory access latency, disk loading latency, and communication latency on each device; $\kappa$ is a constant coefficient; $k$ is the number of rounds to predict one token; ${\bm{w}}^{\prime}$ and ${\bm{n}}^{\prime}$ are the extended vectors of ${\bm{w}}$ and ${\bm{n}}$; ${\bm{z}}$ and ${\bm{z}}^{\text{gpu}}$ are constraint vectors for RAM and VRAM; ${\bm{P}}_{w},{\bm{P}}_{n},{\bm{P}}_{n}^{\text{gpu}}$ are diagonal matricies that activate or deactivate the decision variables.
For the derivation of this definition, please see Appendix A.3. Simply put, constraint (2) ensures that each device handles at least one layer, with the GPU layers not exceeding the total. Constraint (3) enforces all devices are assigned an equal number of windows and all the windows are filled. This is not mandatory in our implementation, but can simplify the problem model. Constraints (4) and (5) ensure that RAM and VRAM usage stay within limits. Table 5 summarizes the key symbols.
To construct ${\bm{a}},{\bm{b}},{\bm{c}},\kappa,{\bm{w}}^{\prime},{\bm{n}}^{\prime},{\bm{z}},{\bm{P}}_{w},{\bm{P}}_{n}$, we should categorize devices into one of the following cases: Case 1 (set $\mathcal{M}_{1}$): macOS with Metal disabled and insufficient RAM; Case 2 (set $\mathcal{M}_{2}$): macOS with Metal enabled and insufficient RAM; Case 3 (set $\mathcal{M}_{3}$): Linux and Android with insufficient RAM; Case 4 (set $\mathcal{M}_{4}$): OS with sufficient RAM or low disk speed. For example, for cases 1–3 where devices should overload, RAM usage must stay above a lower bound, but for case 4 where overloading is not allowed, RAM usage must stay below a upper bound.
Whether a device is overloaded depends on ${\bm{w}}$ and ${\bm{n}}$, e.g., a large ${\bm{w}}_{m}-{\bm{n}}_{m}$ will overload RAM. However, we cannot determine a device’s case before solving the problem, and without knowing the case, we cannot solve the problem. This traps us in a circular dependency. Besides, since $W={\bm{e}}^{\mathrm{T}}\cdot{\bm{w}}$ appears in both the objective and constraints, this LDA model is not a standard integer linear fractional programming (ILFP) problem, making the problem more challenging.
3.3 Our Halda Scheduler
To solve the LDA problem, our core ideas include: (i) transform the NP-hard original problem into a set of standard integer linear programming (ILP) problems by enumerating over all possible $k$, and (ii) search for optimal set assignments $\mathcal{M}_{1}\sim\mathcal{M}_{4}$ by iterative optimization.
1 Initialize layer windows ${\bm{w}}$ proportionally to devices’ memory budgets and the number of GPU layers ${\bm{n}}\leftarrow\mathbf{0}$;
@@ 5       Calculate $W={\bm{e}}^{\mathrm{T}}\cdot{\bm{w}}$ and $k=L/W$;
6       Re-assign devices to sets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$ based on the latest ${\bm{w}},{\bm{n}},k$ and $\mathcal{M}_{4}^{\text{force}}$;
@@
11             Solve the ILP problem (1-5) with fixing $k$ using a ILP solver;
@@ 12             Update best solution $({\bm{w}}^{\star},{\bm{n}}^{\star})$ if the current objective is smaller;
14             Force the device $m_{s}$ from $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$ with lowest disk read speed into $\mathcal{M}_{4}^{\text{force}}$;
@@ 17      Update ${\bm{w}}\leftarrow{\bm{w}}^{*}$ and ${\bm{n}}\leftarrow{\bm{n}}^{*}$;
Transform into standard ILP problems. Given that the number of layers $L$ in typical LLMs is less than 100, the integer $k$ has a limited range of values — at most 11 valid factors for any $k\leq 100$. By enumerating over these factors, we can treat $k$ and $W$ as constants, then the problem becomes:
@@
${\bm{P}}_{w}\cdot{\bm{w}}^{\prime}+{\bm{P}}_{n}\cdot{\bm{n}}^{\prime}+W{\bm{z}}<0,$ (9)
@@
Hence, for each fixed $k$, the objective and constraints boil down to linear functions/inequalities, and the problem becomes an ILP. Then, we can run a standard ILP solver (e.g., HiGHS (Huangfu and Hall, 2018)) to obtain the optimum ${\bm{w}},{\bm{n}}$.
Iterative optimization on set assignments. The problem remains unsolvable because the set assignments $\mathcal{M}_{1}\sim\mathcal{M}_{4}$ are unknown. To resolve the circular dependency, we adopt an iterative optimization approach. Initially, ${\bm{w}}$ is set proportionally based on the memory budget ($d_{m}^{\text{avail}}$ for macOS without Metal and Linux, $d_{m,\text{metal}}^{\text{avail}}$ for macOS with Metal, and $d_{m}^{\text{avail}}+d_{m}^{\text{swapout}}$ for Android), and ${\bm{n}}$ is initialized to $\mathbf{0}$. This gives an initial division of devices into sets $\mathcal{M}_{1}\sim\mathcal{M}_{4}$. Then, we solve the ILP problems to update ${\bm{w}}$ and ${\bm{n}}$, re-assign the devices to their new sets, and repeat this process until the set assignment remain unchanged.
This approach still has defects. The ILP model only constrains the upper bound on the number of layers that can be allocated to the GPU ${\bm{n}}$. In certain cases, such as when the total number of layers $L$ is small and when the initial set assignment falls into a local optimum, the number of layers assigned to the GPU may be insufficient. For example, if a device $m$ is assigned $w_{m}=30$ layers and its GPU can handle up to 20 layers, but due to memory constraints in $\mathcal{M}_{1}\sim\mathcal{M}_{3}$, only 10 layers go to the GPU, leading to underutilization of GPU resources. This issue is caused by the initial set assignment, which allocates too many devices to $\mathcal{M}_{1}\sim\mathcal{M}_{3}$, leaving an insufficient number of layers for GPUs. Therefore, in Algorithm 1, we apply a calibration step: if a GPU is underutilized (its VRAM is not full) while another device is overloaded (i.e., its VRAM is full and has offloaded some layers to the CPU, or it has exceeded the RAM limit), we force the device $m\in\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$ with lagging disk into $\mathcal{M}_{4}$, re-construct the sets $\mathcal{M}_{1}\sim\mathcal{M}_{4}$, and re-solve the problem. This way, we can converge to the optimal set assignment and then determine the optimal setup for ${\bm{w}},{\bm{n}}$ and $k$.
Complexity Analysis. The main loop has two key tasks: set assignment and solving the LDA problem. It runs for $T$ iterations, which in the worst case is $T=O(M)$. Set assignment also takes $O(M)$ time. For LDA, we solve an ILP for each valid factor of $L$, with $K=O(\log L)$ factors in total. Our ILP is usually tiny and sparse, taking $O((2M)^{3.5})$ to solve it. Thus, the total time complexity is $O(T(M+K(2M)^{3.5}))$, which can be solved in polynomial time.
@@ 4 Experiments
We built the experimental platform on common home devices detailed in Table 2. On Mate40Pro (HarmonyOS) and Honor Pad (Android), we run prima.cpp inside a Termux-simulated Linux. These devices were connected to a local Wi-Fi router. By default, we used 4 devices (D1–D4) with a total available RAM+VRAM of 37 GiB (not enough for a Q4K-quantized 70B model). We evaluated Llama models from 8B to 70B (in Q4K format) in terms of token latency, TTFT, and memory pressure, and compared with llama.cpp (Gerganov, 2024), exo (Exo, 2024), and dllama (Tadych, 2024). The results showed the significant advantages of prima.cpp in both speed and memory pressure.
Table 2: Configurations on experimental device
@@
OS MacOS (UMA) Linux Linux Linux (on HarmonyOS) Linux (on Android) MacOS (NUMA)
@@
RAM (available) 2.4 GiB 4.1 GiB 9.7 GiB 1.9 GiB 5.1 GiB 6.8 GiB
Disk Read Speed 0.72 GB/s 2.98 GB/s 3.17 GB/s 1.37 GB/s 2.00 GB/s 0.39 GB/s
GPU Type Apple Metal 3070 2080TI - - -
VRAM (available) - 8 GiB 11 GiB - - -
4.1 Faster Inference on Larger Models
As llama.cpp is an on-device system, we ran it on the most powerful desktop. Meanwhile, we ran exo on devices D1-D3, as they are equipped with Apple/NVIDIA GPUs, and running exo on D4 require root access, which is not permitted. Instead, dllama and prima.cpp used devices D1-D4. Table 3 presents the token latency and TTFT across different model sizes. As of now, exo and dllama do not support Llama models from 14B to 65B, so we put a “-" in the table.
As illustrated in Figure 9(a) (in Appendix A.6), for smaller models ($\leq$14B), the 11 GiB VRAM on the desktop is sufficient for full GPU inference, at this time, llama.cpp is the best choice. However, at 30B, VRAM runs out, forcing layers onto the CPU, thus inference speed becomes limited. As the model size increases to 45B, even the 9.7 GiB of available RAM is exhausted. Since llama.cpp uses mmap to lazily load model weights, the OS frees inactive mmap-ed pages and reloads them as needed, which introduces disk loading latency. At this stage, only a few pages are released, so efficiency loss is small. However, at 60B, more active mmap-ed pages are labeled as inactive earlier and then released, leading to a sharp increase in token latency and TTFT. This indicates that llama.cpp is not well-suited for running larger models on home devices.
@@ 4.1 Faster Inference on Larger Models
For exo and dllama, despite only data for the 8B model is available, their limitations are already evident. Llama 3-8B (Q4K) requires just 5.3 GiB of VRAM, so all 32 layers could fit entirely on D3-GPU. However, exo allocates model layers proportional to each device’s memory. D1 (8 GiB RAM), D2 (8 GiB VRAM), and D3 (11 GiB VRAM) are assigned 9, 10, 13 layers, respectively. While D1 has an Apple Metal GPU, its efficiency is much lower than D3-GPU, making it an efficiency bottleneck. For dllama, inference is performed using tensor parallelism, and the workload is evenly distributed across devices. This forces high- and low-performance devices to handle the same load and causes blocking during all-reduce, leading to limited efficiency.
Prima.cpp addresses these limitations effectively, which uses piped-ring parallelism with prefetching to collaborate multiple devices. Unlike existing systems using solely CPU or GPU, prima.cpp uses both CPU and GPU on each device. It also adapts better to heterogeneous home devices by analyzing CPU/GPU power, available RAM/VRAM, memory management behavior, and disk speed to assign workloads, with the goal to minimize token latency. As shown in Table 3, prima.cpp has significantly lower token latency and TTFT than exo and dllama across all model sizes and outperforms llama.cpp for models larger than 30B. Specifically, compared to llama.cpp, prima.cpp improves token latency by up to 17× and TTFT by up to 8×. Against exo and dllama, it speeds up token latency by 5–8× and TTFT by 12–24×. It also effectively prevents OOM errors.
@@
Table 3: Token latency and TTFT (in millisecond/token) for llama.cpp, exo, dllama, and prima.cpp.
Model llama.cpp exo dllama prima.cpp (w/o halda) prima.cpp (w/o prefetch) prima.cpp
Latency TTFT Latency TTFT Latency TTFT Latency Latency Latency TTFT
Llama 3-8B 15 18 263 960 459 1845 78 53 54 78
Llama 3-14B 20 25 - - - - 131 62 65 134
Llama 1-30B 202 611 - - - - 258 79 72 214
Llama 3-45B 328 712 - - - - 409 263 233 440
Llama 3-60B 7965 8350 - - - - 7053 532 468 990
Llama 1-65B 8807 9662 - - - - 12253 688 569 1770
Llama 3-70B 10120 10806 OOM OOM OOM OOM 20848 755 674 1793
Appendix A.4 shows the results on more hot models, including Qwen 2.5, QwQ and DeepSeek R1. In current implementation, each device is assigned at least one model layer, for example, leading to a 1:1:29:1 split for Llama 3-8B. This restriction is unnecessary and we will remove it in future updates. Then, we will have a 0:0:32:0 split and idle devices removed, making llama.cpp a special case of prima.cpp when serving small models. Appendix A.5 explains why we use D1-D4 instead of D1-D6: more is not always better, and the cluster memory does not need to match the model’s needs. It also shows how prima.cpp can pick a subset of devices to build a best-performing cluster.
4.2 Ablation Study on Prefetching, Halda and Piped-ring Parallelism
Table 3 also includes an ablation study on Halda and prefetching. For a fair comparison, we set prima.cpp (w/o halda) to use exo’s strategy - assigning model layers proportional to device RAM/VRAM, but improves it by using available RAM/VRAM instead of total and offloading overloaded layers from GPU to CPU to prevent OOM. Instead, prima.cpp uses our Halda strategy.
Prefetching. To test prefetching, we disable it in prima.cpp (w/o prefetch) and enable it in prima.cpp. Prefetching has no effect on small models, as they fit entirely in RAM/VRAM without triggering memory reclamation. However, for larger models, memory reclamation triggers frequent page faults as early layers have been released by the OS, but computation starts with them. To mitigate this, prima.cpp prefetches released layers immediately after computation, ensuring they are already in memory before the next round of computation begins to reduce page faults. This overlaps disk loading with other devices’ operations, thus lowering token latency by 9$\sim$17%.
Figure 2: Normalized token latency over $k$.
Halda. A proper layer assignment is crucial for fast inference. For small models, RAM/VRAM on D1-D4 is sufficient, so prima.cpp (w/o halda) and prima.cpp perform similarly. However, prima.cpp (w/o halda) assigns more layers to weaker CPUs/GPUs, while prima.cpp prioritizes powerful GPUs, the latter achieves lower latency. For larger models, prima.cpp (w/o halda) overloads non-GPU devices, triggering memory reclamation and reloading, causing a latency spike. In contrast, prima.cpp avoids memory overload on slow-disk devices through optimized layer assignment, achieving up to 30× speedup for 70B-scale models.
Piped-ring Parallelism. To evaluate it in isolation, we built a CPU cluster with 4 Linux devices, each having 8 cores, 8 GiB RAM, and SSD of 2 GB/s. We tested models from 8B to 72B, and assigned model layers evenly across devices. For example, for 65B-72B models with 80 layers, at $k=1$ ($k$ denotes the rounds to predict one token), layers were split 20:20:20:20; at $k=2$, 10:10:10:10, and so on. The results in Figure 2 show that when total available RAM is insufficient ($>$60B), piped-ring parallelism reduces latency by nearly half. However, at $k=1$, it offers no benefit due to the “prefetch-release" effect. Thus, the layer window size should be set smaller, and piped-ring parallelism becomes essential (see Appendices A.1 and A.2). However, when memory is sufficient ($<$45B), no disk loading occurs, but increasing $k$ slightly raises latency due to overhead from model fragmentation and reduced computational parallelism. To sum up, piped-ring parallelism excels under high disk-loading conditions, but $k=1$ is better when memory is sufficient.
Table 4: Memory pressure for llama.cpp, exo, dllama, and prima.cpp on each device.
Model llama.cpp exo dllama prima.cpp

### v2

@@ 3.2 Layer-to-device Assignment Problem
As mentioned in Table 2, prima.cpp is designed to use multiple backends [^3] Our prototype supports CUDA, Metal, and CPU backends, but with minor adaptation, it can also work with others such as Vulkan and ROCm.. Within a layer window, some layers reside in VRAM and execute on the GPU, while others are offloaded to RAM and executed on the CPU. If the layers assigned to the CPU exceed available RAM, the overflow is further offloaded to external storage and reloaded into RAM via mmap [^4] We pin model layers in VRAM and avoid paging them as in disk-RAM; otherwise, disk-RAM and RAM-VRAM transfers would contend for PCIe (or shared) bandwidth, increase disk reload volume, and amplify disk loading latency, which offsets the gain from GPU acceleration.. This leads to two challenges: How to set the layer window size for each device? Which layers run on the GPU and which on the CPU?
Prior work assumes sufficient aggregated VRAM and proposes heuristic partitioning strategies. For example, Exo (2025) partitions model layers based on memory ratio, so devices with more RAM/VRAM handle more layers. Ye et al. (2024) partition by compute power and migrate layers from OOM devices to those with free memory. These heuristics work on some testbeds but are not always optimal (see Fig. 10 in Appendix A.10), since memory size does not guarantee compute power, and under disk offloading, devices with larger memory but weaker CPUs/GPUs can perform worse than those with less memory but faster CPUs and disks.
To achieve optimal partitioning, TPOT must be quantified through an analytical model. Building such a model is challenging: beyond device heterogeneity, we must account for CPU-GPU coordination, memory contention, OS-specific memory reclamation, disk optimizations, and quantization. For instance, some PCs have dedicated GPUs with separate VRAM; others (e.g., Mac M-series) adopt a UMA architecture where CPU and GPU share memory, and OS reclamation will be more aggressive; some are NUMA systems without GPUs, and they differ in reclamation thresholds across OSs; even on the same device, OS reclamation differs with or without Metal; Linux optimizes sequential disk reads, making reloads faster; quantization influences compute latency, memory access, disk loading, and RAM/VRAM constraints, ultimately shaping the partitioning strategy.
After extensive performance analysis in Appendix A.3 and preliminary experiments, we formalize this as the layer-to-device assignment (LDA) problem as follows.
Definition 1 (Layer-to-device assignment).
Assume there are $M$ devices, $w_{m}$ is the layer window size on device $d_{m}$ and $n_{m}$ is the number of GPU layers within $w_{m}$. Let the decision variables be ${\bm{w}}^{\mathrm{T}}=[w_{1},w_{2},\cdots,w_{M}]$ and ${\bm{n}}^{\mathrm{T}}=[n_{1},n_{2},\cdots,n_{M}]$. Our objective is to find the optimal $\bm{w}$} and $\bm{n}$} that minimizes the TPOT:
@@
${\bm{P}}_{w}\cdot{\bm{w}}^{\prime}+{\bm{P}}_{n}\cdot{\bm{n}}^{\prime}+{\bm{e}}^{\mathrm{T}}\cdot{\bm{w}}\cdot{\bm{z}}\leq 0,$ (4)
@@
where $L$ is the number of model layers; ${\bm{a}},{\bm{b}},{\bm{c}}$ are latency coefficient vectors determined by compute latency, memory access latency, disk loading latency, and communication latency on each device; $\kappa$ is a constant latency offset; $k$ is the number of rounds to predict one token; $\bm{w}$$^{\prime}$ and $\bm{n}$$^{\prime}$ are the extended vectors of $\bm{w}$} and $\bm{n}$}; $\bm{z}$} and ${\bm{z}}^{\text{gpu}}$ are constraint vectors for RAM and VRAM; ${\bm{P}}_{w},{\bm{P}}_{n},{\bm{P}}_{n}^{\text{gpu}}$ are diagonal matrices that activate or deactivate the decision variables; and $\bm{e}$} is an all-ones vector.
Constraint (2) ensures the layers do not exceed the limit. Constraint (3) enforces that all devices are assigned an equal number of windows and all windows are filled [^5] This is not mandatory in our implementation, but it can simplify the problem model.. Constraints (4) and (5) ensure that RAM and VRAM usage stay within limits. Table 6 summarizes the key symbols. To construct ${\bm{a}},{\bm{b}},{\bm{c}},\kappa,{\bm{w}}^{\prime},{\bm{n}}^{\prime},{\bm{z}},{\bm{P}}_{w},{\bm{P}}_{n}$, we categorize devices into four sets: (a) Set $\mathcal{M}_{1}$: macOS with Metal disabled and insufficient RAM; (b) Set $\mathcal{M}_{2}$: macOS with Metal enabled and insufficient RAM; (c) Set $\mathcal{M}_{3}$: Linux and Android with insufficient RAM; (d) Set $\mathcal{M}_{4}$: devices with sufficient RAM or slow disk. For Sets $\mathcal{M}_{1}$-$\mathcal{M}_{3}$ where devices should overload, RAM usage should stay above available RAM; for Set $\mathcal{M}_{4}$, where overloading is not allowed, RAM usage should stay below available RAM. We can extend other OSs to these sets, or create a new set and adjust variable dimensions.
This problem is an NP-hard integer linear fractional program (ILFP): both the numerator and denominator of the objective are linear in the decision variables, and all constraints are linear inequalities. Whether a device is overloaded depends on $\bm{w}$} and $\bm{n}$}, e.g., a large ${\bm{w}}_{m}-{\bm{n}}_{m}$ will overload RAM. However, we cannot determine a device’s set before solving the LDA problem, and without the set assignment, we cannot solve the LDA problem. This traps us in a circular dependency.
3.3 Halda: Automatic Layer Partitioning and Device Selection
To solve this non-standard, NP-hard LDA problem, our core ideas include: (i) transform the original problem into a set of standard integer linear programs (ILPs) by enumerating over all valid $k$ that divide $L$, and (ii) search for optimal set assignments $\mathcal{M}_{1}$-$\mathcal{M}_{4}$ by iterative optimization.
1Initialize layer windows $\bm{w}$} proportionally to devices’ memory budgets, and GPU layers ${\bm{n}}\leftarrow\mathbf{0}$;
@@ 5   Calculate $W={\bm{e}}^{\mathrm{T}}\cdot{\bm{w}}$ and $k=L/W$;
6   Reassign devices to sets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$ based on the latest ${\bm{w}},{\bm{n}},k$ and $\mathcal{M}_{4}^{\text{force}}$;
@@
11      Solve the ILP for fixed $k$ using a solver;
@@ 12      Update best solution $({\bm{w}}^{\star},{\bm{n}}^{\star})$ if the current objective is smaller;
14      Force the device $m_{s}$ from $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ with slowest disk reads into $\mathcal{M}_{4}^{\text{force}}$;
@@ 17  Update ${\bm{w}}\leftarrow{\bm{w}}^{*}$ and ${\bm{n}}\leftarrow{\bm{n}}^{*}$;
Transform into standard ILPs. Given that the number of layers $L$ in LLMs is typically less than 100, the integer $k$ has a limited range of values: at most 11 valid factors for any $L\leq 100$. By enumerating these factors, we treat $k$ and $W$ as constants, and the problem becomes:
@@
${\bm{P}}_{w}\cdot{\bm{w}}^{\prime}+{\bm{P}}_{n}\cdot{\bm{n}}^{\prime}+W{\bm{z}}\leq 0,$ (9)
@@
Hence, for each fixed $k$, the objective and constraints boil down to linear functions/inequalities, and the problem becomes an ILP. Then, we can run a standard ILP solver (e.g., HiGHS (Huangfu & Hall, 2018)) to obtain the optimal ${\bm{w}},{\bm{n}}$.
[…]
However, this approach still has defects. For example, if a device $m\in\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ is assigned 30 layers, although the GPU can host 20 layers, memory constraints may end up assigning only 10 to the GPU and 20 to the CPU, underutilizing the GPU. This issue arises from an improper set initialization: if the device was initialized in $\mathcal{M}_{4}$, its GPU could be fully utilized. Therefore, Algorithm 1 includes a calibration step: if a GPU is underutilized (VRAM not full) while another device is overloaded (VRAM full with layers offloaded to CPU, or RAM exceeded), we move the device with slowest disk reads from the set $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ into $\mathcal{M}_{4}^{\text{force}}$ ($\mathcal{M}_{4}^{\text{force}}\subset\mathcal{M}_{4}$) and solve again. This ensures convergence to the optimal set assignment and the corresponding optimal values of $\bm{w}$}, $\bm{n}$}, and $k$.
Complexity analysis. The main loop alternates between set assignment and LDA solving, repeated for $T$ iterations with $T=O(M)$ in the worst case. The set assignment takes $O(M)$. For LDA, we solve a tiny ILP for each valid factor of $L$, giving $K=O(\log L)$ factors in total. Although ILPs are NP-hard, our instances are small and sparse, allowing modern solvers to finish quickly. Empirically, on 4-32 devices, the global scheduling latency is 10-12 ms. The overhead of rerunning Halda is negligible, and since no cross-device layer migration is required, workloads can be repartitioned whenever the task queue is empty to adapt to environmental changes.
Device selection. Weak devices assigned only one layer are removed, since Halda indicates that excluding them would improve speed. Appendix A.7 illustrates this selection process. For small models, Halda prefers to allocate all layers to one most powerful device, reducing prima.cpp to llama.cpp. However, if removing a device would break the communication ring (e.g., it is the only reachable hop due to network policy), Halda keeps it in the ring as a relay node but assigns it no workload. This design frees users from device selection: we only need to discover more available devices, and Halda will select the optimal subset for us.
@@ 4 Experiments
We implement prima.cpp with 20K lines of code based on llama.cpp and build a real testbed of heterogeneous, low-end home devices (see Table 3). These devices connect via a local Wi-Fi router, with inter-device bandwidth ranging from 320-610 Mbps and link latency from 3-7 ms. By default, four devices (D1-D4) with aggregated RAM+VRAM of 37 GiB (insufficient for a Q4K-quantized 70B model) are used. We evaluated Llama models from 8B to 70B (Q4K) in terms of time-per-output-token (TPOT), time-to-first-token (TTFT), and memory pressure. We select llama.cpp, exo, and dllama as baselines, as they are the most popular open-source on-device (distributed) LLM inference systems [^7] At submission, llama.cpp had 87K stars, exo 31K stars, and dllama 2.7K stars. We also tried to deploy vLLM and SGLang, but as server-oriented systems, they failed to run on our home cluster.. Since llama.cpp is a standalone on-device system, we run it on D3, which offers the largest RAM/VRAM and highest decoding efficiency [^8] At small batch sizes, decoding is bandwidth-bound, and 2080TI offers higher memory bandwidth than 3070.. Exo is a PP system, which we run on D1-D3, but D4 is unavailable because it requires root access. In contrast, dllama and prima.cpp run on D1-D4.
Table 3: Testbed configuration. Termux is used on D4 and D5.
@@
OS macOS Linux Linux HarmonyOS Android macOS
@@
RAM (avail.) 2.4 GiB 4.1 GiB 9.7 GiB 1.9 GiB 5.1 GiB 6.8 GiB
Disk Read 0.7 GB/s 3.0 GB/s 3.0 GB/s 1.4 GB/s 2.0 GB/s 0.4 GB/s
GPU Type Apple Silicon 3070 2080TI - - -
VRAM (avail.) - 8 GiB 11 GiB - - -
4.1 Fast Inference on Large Models
[…]
For small models (<14B), Halda finds that D3 can hold the entire model in VRAM, so it removes the other devices and runs only on D3. This reduces prima.cpp to the single-device case (equivalent to llama.cpp), with both achieving the same TPOT and TTFT. At 30B, as shown in Fig. 9(a) in Appendix A.8, D3-GPU runs out of VRAM and forces layers onto the CPU, causing llama.cpp to deteriorate rapidly. In contrast, Fig. 9(d) shows that Halda offloads the overflow from D3-GPU to D2-GPU, keeping TPOT and TTFT stable. At 45B, both RAM and VRAM on D3 are exhausted, and mmap in llama.cpp begins frequent reloads, incurring disk latency. At this stage, only a few pages are reloaded, so the efficiency loss is minor. At 60B, high memory stress causes more active pages to be evicted earlier, sharply increasing TPOT and TTFT. This indicates that llama.cpp is ill-suited for large models on consumer-grade devices. In contrast, Halda balances workloads across devices according to their capabilities. Even when disk offloading is required, it assigns workloads to devices with strong CPU and fast disk, while PRP with prefetching hides disk latency. As a result, prima.cpp maintains stable TPOT and TTFT growth and achieves sub-second TPOT even at 70B. To the best of our knowledge, prima.cpp is the first system to achieve this speed on real consumer-grade, low-end devices [^10] We exclude sparse inference systems, as they do not meet the conditions in Table 1: they require dedicated hardware such as NPUs, support only sparse models, and may degrade output quality..
@@ 4.1 Fast Inference on Large Models
For exo and dllama, despite only 8B and 70B being available, their limitations are already evident. Llama 3-8B (Q4K) requires 5.3 GiB of VRAM, so the entire model could fit into D3-GPU. However, exo allocates layers proportional to each device’s memory. D1 (8 GiB RAM), D2 (8 GiB VRAM), and D3 (11 GiB VRAM) are assigned 9, 10, 13 layers, respectively. While D1 has an Apple Silicon GPU, its efficiency is much lower than the 2080TI on D3-GPU, making it an efficiency bottleneck.
For dllama, inference is based on TP. For the 8B model, 64 all-reduces per token incur at least 192 ms of delay under Wi-Fi latency (3-7 ms), making its measured TPOT (459 ms/token) far above prima.cpp (15 ms/token). At 70B, dllama runs out of memory, but our independent tests show $\sim$150 ms per all-reduce, so 160 all-reduces add 24 s/token, yielding $\sim$30 s/token compared to prima.cpp’s 674 ms/token (442 ms/token with speculative decoding). These results highlight that Halda and PRP are better suited for heterogeneous, high-latency home environments than heuristic partitioning or TP.
@@
Table 4: TPOT (ms/token) and TTFT (ms) on Llama models for llama.cpp, exo, dllama, prima.cpp.
Size llama.cpp exo dllama prima.cpp (w/o halda) prima.cpp (w/o prefetch) prima.cpp
TPOT TTFT TPOT TTFT TPOT TTFT TPOT TPOT TPOT TTFT
8B 15 18 263 960 459 1845 78 15 15 18
14B 20 25 - - - - 131 20 20 25
30B 202 611 - - - - 258 79 72 214
45B 328 712 - - - - 409 263 233 440
60B 7965 8350 - - - - 7053 532 468 990
65B 8807 9662 - - - - 12253 688 569 1770
70B 10120 10806 OOM OOM OOM OOM 20848 755 674 1793
Similar results are observed on another testbed in Appendix A.4 and a homogeneous testbed in Appendix A.5, confirming that prima.cpp’s speedups generalize across diverse testbeds. Appendix A.6 shows the results on more models, including Qwen 2.5, QwQ, and DeepSeek R1. Appendix A.7 explains the underlying rationale for device selection: more is not always better, and aggregated memory need not match the model’s needs. It also shows how Halda selects a subset of devices to build a best-performing cluster. Appendix A.9 shows that with speculative decoding, 32B inference achieves 26 tokens/s, offering both the speed and intelligence needed for LLM agents.
4.2 Ablation Study on Prefetching, Halda, and Pipelined-Ring Parallelism
Table 4 also presents ablations on Halda and prefetching. Exo is also a PP system, so we migrate exo’s layer partitioning (based on memory ratio) to prima.cpp (w/o halda), making it an exo variant. Unlike exo, this variant uses available instead of total memory, adds a broader range of models, adds cross-platform quantization and prefetching, and supports CPU/disk offloading to prevent OOMs.
Prefetching. To evaluate prefetching, we compare prima.cpp with and without it. It has little effect on small models that fit in RAM/VRAM, but for larger models, evicted layers cause frequent page faults and reloads. Without prefetching, all reloads are triggered on demand by page faults. With prefetching, upcoming layers are loaded in advance and their latencies overlapped, reducing page fault reloads and lowering TPOT by 9-17%.
[…]
Figure 2: TPOT of PRP under different rounds $k$.
PRP with multiple per-token rounds. To evaluate PRP in isolation, we built a CPU cluster with 4 devices, each with 8 cores, 8 GiB RAM, and an SSD of 2 GB/s. We tested models from 8-72B, and assigned layers evenly across devices. For example, for models with 80 layers, at the per-token round $k=1$, layers were split 20:20:20:20; at $k=2$, 10:10:10:10, and so on. With large models (>60B) and $k=1$, PRP degrades to PP and provides no benefit due to the prefetch-release conflict (see Appendix A.1). In contrast, with $k\geq 2$, Fig. 2 shows that PRP halves TPOT by eliminating the prefetch-release conflict and allowing finer-grained overlap of disk loading (see Appendix A.2). With small models (<45B) and sufficient memory, the model fits in the cluster without disk offloading. Then, increasing $k$ has minimal impact on TPOT, aside from minor overhead from additional communication and kernel launches. This shows that PRP is more effective than PP for 70B-scale models.
Table 5: Memory pressure for each device on Llama models.
Size llama.cpp exo dllama prima.cpp